OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 minus 1 over x minus 3 equals 1 over 4x minus 12 ? @midhun.madhu1987

OpenStudy (campbell_st):

so the equation looks like \[\frac{x}{x^2 - 9} + \frac{1}{x -3} = \frac{1}{4x - 12}\] factor the denominator of the 1st fraction and the last and you get \[\frac{x}{(x + 3)(x -3)} + \frac{1}{(x -3)} = \frac {1}{4(x - 3)}\]

OpenStudy (campbell_st):

next multiply every fraction by (x - 3) and you get \[\frac{x}{x + 3} + 1 = \frac{1}{4}\] subtract 1 from both sides \[\frac{x}{(x + 3)} = - \frac{3}{4}\] cross multiply then which gives \[4x = -3(x + 3)\] now solve for x